how do you find the domain and range
3. (10 points) Find the inverse of the given function using these steps 1. Replacef(x) withy. 2. Interchange x and y. 3. Solve for y. 4·Rep lacey with广1(x). X- +/3对 Domain of fox): p Domain off-1(x): し aotuta 9- 3 f(x) = x-3 Uncti INVERSE Domaim Range of f x): -多す
Answer #1

\\f(x) = \frac{2}{x-3}\\ \\1.\ y =\frac{2}{x-3}\\ \\2.\ x = \frac{2}{y-3}\\ \\3.\ y-3 = \frac{2}{x}\\ \\y = \frac{2}{x}+3 = \frac{3x+2}{x}\\ \\4.\ f^{-1}x = \frac{3x+2}{x}\\\

please upvote the answer.